l Bausher Investment Company is fond of investing in equity shares listed in Muscat Securities Market (MSM). During COVID-19 Pandemic, most of the company’s share prices have fallen down. Al Bausher Investment Company has sold a stock of 100 shares each day at RO 12 per share on first day, RO 10 on the second day and RO 9 on the third day. The company has purchased 50 shares on the first day, 80 shares on the second day and 100 shares on the third' day of Gulf Construction company using the amount company collected by selling the shares .
a) What mean price per share did Al Bausher Investment Company pay to buy the shares?
b) One week after the shares are bought, there is 10% spike in the share prices of Gulf Construction Company. Should Al Bausher sell the shares of Gulf Construction Company? Discuss analytically.   
Answer #1

Selling of MSM:

Collection of Day1 : 100*12=1200

Collection of Day2 : 100*10 =1000

Collection of Day 3: 100*9 =900

Buying of Gulf Construction:

Day1 (50 shares) : Avg, Buying Price: (1200 / 50) =24

Day2 (80 shares) : Avg. Buying Price: (1000 / 80) =12.5

Day 3 (100 shares) : Avg. Buying Price: (900 / 100) = 9

a)

Therefore mean Price =  \frac{(24*50)+(12.5*80)+(9*100)}{50+80+100}

=\frac{(1200)+(1000)+(900)}{230}

=\frac{3100}{230}

= 13.4783

b)

(Since it is not mentioned in the question, so we are assuming the price remained constant i. e. 9 per share before the spike)

Now, there is 10% Spike in Prices

i.e. Current Price = (9 + 10% of 9)

= 9.9

Even after the spike, the price is below the mean price mean price i.e. Al Bausher is still in loss.

And to reach the mean price of buying, the share price need a spike of about 13.6%.

Since the share has already seen a downfall and is showing a spike, Al Bausher should not sell the shares and hold them.

Kindly UpVote. It Helps. Thank You